OpenStudy (anonymous):

Verify: cot[x-(pi/2)]=-tan(x)

mathslover (mathslover):

\(\cot \theta = \cfrac{\cos \theta}{\sin \theta}\) so, here u have : \(\cot (x - \pi /2) = \cfrac{\cos (x - \pi /2 ) }{\sin (x -\pi /2) }\)